OpenStudy (anonymous):

What is the value of a in the following equation: 2a = -a + 9? 3 -6 -3 none of these

OpenStudy (awesome781):

none of these

OpenStudy (phi):

2a = -a + 9 add +a to both sides what do you get ?

OpenStudy (awesome781):

or u could just plug in the numbers and see if both sides equal each other....more easier..

OpenStudy (awesome781):

That is what I did....

OpenStudy (anonymous):

@Awesome781 :)

OpenStudy (phi):

2a = -a + 9 add +a to both sides 2a+a = -a + a + 9 on the left side , what is 2 a's plus 1 more a? on the right side what is 1 a take away 1 a leaving just the 9 ?